OpenStudy (anonymous):

solve 4+3^5x=8 for x.

OpenStudy (anonymous):

Is it ? \[4+3^{5 x}=8 \] @andria_elizabeth

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Subtract 4 from both the sides, what would you get?

OpenStudy (anonymous):

3 ^5x=4

OpenStudy (anonymous):

Do you know logarithms?

OpenStudy (anonymous):

a little bit but thats what im confused on, can you please help me out?

OpenStudy (anonymous):

Yeah, Let's take log both the sides, base is 10 \[\log 3^{5x}=\log 4\] Now log has a property \[\log a^b=b\log a\] Let's use this here \[\log 3^{5x} =\log 4\] \[ 5x \log 3= \log 4\] Do you get this?

OpenStudy (anonymous):

why is it log 4?

OpenStudy (anonymous):

is it proabably good for me to memorize the properties?

OpenStudy (anonymous):

You have 4 on the right side, we applied log to both the sides. That's how we get \[5x\times \log 3=\log 4\] Yeah, you should memorize all logarithmic properties.

OpenStudy (anonymous):

@andria_elizabeth Now we should refer the log tables, to find the value of log 4 and log 3. Both are to base 10

OpenStudy (anonymous):

okay so even if it didnt have log before we still put it? and how do we find the value tho?

OpenStudy (anonymous):

Let's see the definition of log \[\log_b a =c\] \[=> a=b^c\] So if you have \[\log_{10} 100=2\] This implies \[100=10^2\] which is true. Do you get it? We could have any base bu it should be >0 and not equal to 1

OpenStudy (anonymous):

yeah, i understand that part

OpenStudy (anonymous):

so here we have log to base 10 If you refer log tables or use calulator \[\log 3=0.477\] \[\log 4=0.6020\] Now we have \[3x\times 0.477=0.602\] Can you find x from here?

OpenStudy (anonymous):

i just take .477 away from .602?

OpenStudy (anonymous):

Nope, you have to divide 0.602 by 0.477

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

\[3x=\frac{0.602}{0.477}\]

OpenStudy (anonymous):

once we divide that then we divide again by 3x?

OpenStudy (anonymous):

yes. Just 3 you need to find x

OpenStudy (anonymous):

oh okay, thanks

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

:]